# Multi-Scale Data Assimilation in Turbulent Models

### v1

@@ IV Numerical Setup
In this study, we work under the perfect-model setting, implying that the Sabra shell model used as ground truth, and about which synthetic data are simulated, is the same as the model used to assimilate such data. The prediction error arises only from imperfect initial conditions. The model parameters are set to $(a,b,c)=(1,-0.5,-0.5)$, with the number of shells $N=20$, viscosity $\nu=10^{-6}$, and time step $dt=10^{-5}$. A constant forcing term, $f_{n}=(1+i)\delta_{n0}$, is applied exclusively to the first shell. The selection of these parameters ensures the stability of the numerical solver as detailed in Appendix A.
Since both the ensemble and the ground-truth solution evolve under the same governing equations, the setup constitutes a twin experiment in DA jargon. Initial conditions’ errors are (generally) amplified because of the chaotic nature of the underlying dynamics. To ensure that the initial conditions for the ground truth, $\bm{u}$, are sampled from the system attractor, that is they are statistically consistent with the system dynamics, a preliminary spin-up integration is carried out over a sufficiently long time interval, over many turnover times $\tau_{0}$ of the first-forced shell (Eq. (5)). This guarantees that we assimilate data from a regime characterized by well-established statistical properties. For EnKF, a larger initial spread allows the ensemble to explore a wider region of the state space, making it more compatible with the first observation to be assimilated and thereby improving filter stability. This also helps mitigating EnKF-specific instabilities and can reduce reliance on artificial inflation. To achieve this, the ensemble is freely evolved for $10\tau_{0}$ before assimilation begins. During this spin-up phase, chaotic dynamics naturally increases the ensemble spread, although it may also reduce the individual member diversity by making them align along the most unstable modes. Since chaotic systems rapidly lose memory of their initial conditions, the specific form of the initialization becomes irrelevant and can therefore be chosen as:
$\hat{u}_{n}^{(j)}(t_{0})=\sqrt{|u_{n}(t_{0})|^{2}}\left[\cos\left(\phi_{n}^{(j)}\right)+i\sin\left(\phi_{n}^{(j)}\right)\right]\ .$ (12)
This approach ensures dynamical balance within the ensemble and promotes the emergence of physically meaningful multivariate correlations [11].
@@
[…]
@@
$\frac{|\max_{N_{\textit{exp}}}(\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}})+\min_{N_{\textit{exp}}}(\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}})|}{2}$ (17)
$\pm$ $\frac{|\max_{N_{\textit{exp}}}(\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}})-\min_{N_{\textit{exp}}}(\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}})|}{2}\,,$
@@
Figure 3: (a) Real parts of the velocity components for shells $n=1$, $6$, and $13$ from an experiment where $u_{6}$, $u_{7}$, and $u_{8}$ are measured with observation interval $\Delta t_{\text{obs}}=\tau_{15}=0.002\tau_{0}$. The figure displays the full ensemble (light blue), its mean (solid blue line), and the ground truth (solid black line), over a total experiment duration of $20\tau_{0}$. (b) Probability density function (PDF) of the normalized real-part error, $\frac{\Re(u_{n}-\tilde{u}_{n})^{2}}{\sqrt{\langle\Re(u_{n})^{2}\rangle_{T}}\sqrt{\langle\Re(\tilde{u}_{n})^{2}\rangle_{T}}}$, and (c) PDF of the phase difference, $\theta_{n}-\tilde{\theta}_{n}$, shown for all ensemble members (shaded area) and their mean (blue line) for the selected shells. The PDFs in (b) and (c) are computed by collecting data only after the transient phase—i.e., after saturation is reached (as indicated by the red vertical line in panel (a))—and up to the end of the $20\tau_{0}$ experiment. The light-green shaded zone highlights the time window $12\leq t/\tau_{0}\leq 14$ detailed in Fig. 4(a) and (b).
Figure 4: (a) Ground truth energy spectrum $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$, EnKF estimate $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$, and $\ell^{2}$-error $\mathcal{E}_{n}=\langle|u_{n}-\tilde{u}_{n}|^{2}\rangle_{T,L}$. The grey shaded area highlights the observed shells, while the vertical solid line marks the scale for which the turnover time matches the observation interval, $\tau_{15}=\Delta t_{\textit{obs}}$. (b) Normalized error $\mathcal{E}_{n}/\sqrt{E_{n}\tilde{E}_{n}}$, with the dashed line indicating the scale beyond which statistical consistency ($E_{n}=\tilde{E}_{n}$) is maintained, but the EnKF no longer achieves effective reconstruction. All quantities in panels (a) and (b), including error bars, are averaged over $N_{\text{exp}}=16$ independent experiments, as described in Section IV. (c) Normalized energy evolution for inertial-range scales during the stationary synchronization regime ($12\leq t/\tau_{0}\leq 14$; see light-green shaded region in Fig. 3(a)). Shown are the ground truth values $|u_{n}|^{2}k_{n}^{2/3}$ (left), one randomly selected posterior ensemble member $|\tilde{u}_{n}|^{2}k_{n}^{2/3}$ (middle), and their difference (right). Red vertical lines indicate the measured shells ($n=6$, $7$, and $8$). (d) Cumulative instantaneous energy flux evolution over the same stationary synchronization regime, shown for the ground truth (left), one randomly selected posterior ensemble member (middle), and their difference (right).
@@ V Ensemble Kalman Filter Results
In our parameter setting, the turnover time of the slowest scale is $\tau_{0}\approx 0.5$, while at the fastest dissipative cutoff scale, it decreases to approximately $\tau_{15}=0.002\tau_{0}$. Sampling at these respective rates (that is, using $\Delta t_{\textit{obs}}=\tau_{n}$) provides discrete or (quasi-) continuous-time information on inertial range dynamics.
@@
In the first set of experiments, observations are taken from shells $n=6,7,8$, using an observation interval of $\Delta t_{\textit{obs}}=\tau_{15}$ and an ensemble size of $L=1000$ members. Figure 3(a) provides a qualitative overview of the assimilation performance, showing the temporal evolution of the EnKF-based estimates for shells $n=1$, $6$, and $13$. The ensemble estimates are continuously updated across all scales, rapidly synchronizing the directly observed $u_{6}$ (middle panel), whose ensemble spread quickly converges within the measurement error bounds. Dynamical correlations also enable full synchronization of distant, unobserved shells such as fast $u_{13}$ and $u_{1}$, albeit with a slight delay. With real-part synchronization already established (Fig. 3(b)), the narrow PDF of $\theta-\tilde{\theta}$ (Fig. 3(c)) completes the reconstruction picture by confirming phase alignment, which in turn implies synchronization of the imaginary components as well.
@@
Figure 6: (a) Ground truth energy $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$ and estimated energy $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$ for different assimilation windows $\Delta t_{\textit{obs}}$. The inset shows the corrispondence with scale turnover times $\tau_{n}$. The grey shaded area highlights the measured scales 6, 7, and 8. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(14)) normalized with energies, for various observation frequencies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. (c) Flux-based error $\mathcal{E}^{\Pi}_{n}$ (Eq.(15)) normalized (see Evaluation Metrics for details on normalization). All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.
@@
In Fig. 4, we summarize the main quantitative results of this first DA experiment. First, note that the EnKF accurately reproduces the statistical properties of the ground truth, as evidenced in Fig. 4(a) by the perfect overlap between the ground-truth energy spectrum $E_{n}$ and its EnKF-based estimate $\tilde{E}_{n}$. More interestingly, the instantaneous reconstruction is also reasonably accurate. This is assessed by comparing the energy spectrum $E_{n}$ with the reconstruction $\ell^{2}$-error $\mathcal{E}_{n}$, as defined in Eq. (14). As shown, the reconstruction is nearly perfect on the observed shells (gray region), and remains highly accurate for all slower shells ($n<6$). For faster scales, good reconstruction is achieved up to $n=16$, where $\mathcal{E}_{n}<E_{n}$, while at higher shell numbers ($n\geq 17$), the reconstruction becomes purely statistical with $\mathcal{E}_{n}\approx 2E_{n}$.
@@
The quality of the instantaneous reconstruction is further quantified in Fig. 4(b), where we plot the dimensionless $\ell^{2}$-error $\mathcal{E}_{n}/\sqrt{E_{n}\tilde{E}_{n}}$, normalized by the energy spectra of the ground truth and assimilated fields (Eq. (16)). Synchronization with ground truth is nearly perfect up to the fastest observed shell ($n=8$), followed by a partially synchronized range ($9\leq n\leq 16$). Panels (c) and (d) provide a qualitative overview of the time-dependent performance of data assimilation (DA), showing both the normalized energy, $k_{n}^{2/3}|u_{n}|^{2}$, and the energy flux, $\Pi_{n}$ (see Eq. (6)), up to the last inertial shell ($n=13$), over a time window encompassing several energy bursts propagating from large to small scales. The left, central, and right panels, respectively, display the ground-truth quantity, a randomly selected ensemble member, and their pointwise difference. A key observation is that, although the presence of dark halos for $1\leq n\leq 8$ (third column of Panel (c)) confirms the accurate instantaneous tracking of both measured and large-scale dynamics, the most pronounced desynchronization events—evident in the same panel and even more so in the highly intermittent flux shown in Panel (d)—are strongly correlated with intense energy bursts, predicted with high temporal fidelity but limited magnitude accuracy.
@@ V.1 Performance Against Measurements Frequency
The magnitude of the error bars also reflects the sensitivity to the inflation factor (as discussed in Appendix B) used for stabilization. For example, the curve corresponding to $\tau_{7}$ in Fig. 6(b) exhibits a pronounced increase in error at scales beyond those directly observed, signaling that assimilation was less effective and stable in those cases. This pattern aligns with the theoretical insights of Gottwald and Majda [17], who showed that the stability of the filter critically depends on the observation interval $\Delta t_{\textit{obs}}$. Specifically, when observations are too frequent, the forecast step is too short for the model dynamics to evolve to overly constrained and potentially unrealistic states. Conversely, when observations are too sparse, the ensemble can spread sufficiently to capture the attractor structure. In other words, while filter divergence is unlikely at either extreme (very small or very large $\Delta t_{\textit{obs}}$) there exists an intermediate regime where the ensemble may be neither sufficiently constrained by observations nor sufficiently dispersed to maintain dynamical consistency, resulting in degraded performance and potential divergence.
The total errors are summarized in Table 1 by summing the errors from the first scale (excluding the forced one) up to the dissipative cutoff at scale 15: $\mathcal{E}=\sum_{n=1}^{15}\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}}$ and $\mathcal{E}^{\Pi}=\sum_{n=1}^{15}\mathcal{E}_{n}^{\Pi}/\sqrt{\tilde{\Delta}_{n}\Delta_{n}}$. A similar procedure is applied to the error bars to obtain the total uncertainties $\Delta\mathcal{E}$ and $\Delta\mathcal{E}^{\Pi}$. Furthermore, the last column of Table 1 reports the percentage of independent assimilation experiments (out of $N{{}_{\text{exp}}}=16$) that required the correct adjustment of the inflation strength factor $\lambda$, as defined in Eq. (33).
@@
6, 7, 8 ±$\mathcal{E}\pm\Delta\mathcal{E}$ $\mathcal{E}_{\Pi}\pm\Delta\mathcal{E}_{\Pi}$ $N_{\text{inflation}}$
$\Delta t_{\text{obs}}=\tau_{15}=0.002\tau_{0}$ $4.85\pm 0.26$ $5.11\pm 0.85$ $6\%$
$\Delta t_{\text{obs}}=\tau_{9}=0.02\tau_{0}$ $7.75\pm 0.17$ $8.13\pm 0.98$ $6\%$
$\Delta t_{\text{obs}}=\tau_{8}=0.04\tau_{0}$ $9.10\pm 0.69$ $10.07\pm 3.21$ $38\%$
$\Delta t_{\text{obs}}=\tau_{7}=0.06\tau_{0}$ $14.68\pm 3.90$ $17.76\pm 8.35$ $50\%$
$\Delta t_{\text{obs}}=\tau_{6}=0.1\tau_{0}$ $22.44\pm 1.14$ $27.89\pm 5.20$ $0\%$
$\Delta t_{\text{obs}}=\tau_{4}=0.2\tau_{0}$ $23.73\pm 0.50$ $29.49\pm 4.23$ $0\%$
Table 1: The total error is computed by summing the normalized errors from the first scale $n=1$ (excluding the forced one) up to the dissipative cutoff at scale $n=15$: $\mathcal{E}=\sum_{n=1}^{15}\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}}$ and $\mathcal{E}_{\Pi}=\sum_{n=1}^{15}\mathcal{E}_{n}^{\Pi}/\sqrt{\tilde{\Delta}_{n}\Delta_{n}}$. A similar summation is applied to the error bars to obtain $\Delta\mathcal{E}$ and $\Delta\mathcal{E}^{\Pi}$. Details on the percentage of experiments (out of $N_{\text{exp}}=16$) that required proper tuning of the inflation strength factor are also provided.

### v2

@@ IV Numerical Setup
In this study, we work under the perfect-model setting, implying that the Sabra shell model used as ground truth, and from which synthetic data are extracted, is the same as the model used to assimilate such data. The prediction error arises only from imperfect initial conditions. The model parameters are set to (a,b,c)=(1,0.5,0.5)(a,b,c)=(1,-0.5,-0.5), with the number of shells $N=20$, viscosity $\nu=10^{-6}$, and time step $dt=10^{-5}$. A constant forcing term, $f_{n}=(1+i)\delta_{n0}$ ($\delta_{n0}$ is the Kronecker symbol), is applied exclusively to the first shell. The selection of these parameters ensures the stability of the numerical solver as detailed in Appendix A.
Since both the ensemble and the ground-truth solution evolve under the same governing equations, the setup constitutes a twin experiment in DA jargon. Initial conditions’ errors are (generally) amplified because of the chaotic nature of the underlying dynamics. To ensure that the initial conditions for the ground truth, $\bm{u}$, are sampled from the system attractor, that is they are statistically consistent with the system dynamics, a preliminary spin-up integration is carried out over a sufficiently long time interval, over many turnover times $\tau_{0}$ of the first-forced shell (Eq. (5)). This guarantees that we assimilate data from a regime characterized by well-established statistical properties. For EnKF, a larger initial spread allows the ensemble to explore a wider region of the state space, making it more compatible with the first observation to be assimilated and thereby improving filter stability. This also helps mitigating EnKF-specific instabilities and can reduce the need for artificial inflation. To achieve this, the ensemble is freely evolved for $10\tau_{0}$ before assimilation begins. During this spin-up phase, chaotic dynamics naturally increases the ensemble spread, although it may also reduce the individual member diversity by making them align along the most unstable modes. Since chaotic systems rapidly lose memory of their initial conditions, the specific form of the initialization becomes irrelevant and can therefore be chosen as:
$\hat{u}_{n}^{(j)}(t_{0})=\sqrt{|u_{n}(t_{0})|^{2}}\left[\cos\left(\phi_{n}^{(j)}\right)+i\sin\left(\phi_{n}^{(j)}\right)\right]\ ,$ (14)
where the phases $\phi_{n}^{(j)}$ are independent random variables uniformly chosen in $0,2\pi]$. This approach ensures dynamical balance within the ensemble and promotes the emergence of physically meaningful multivariate correlations [11].
@@
with $m,n=0,\ldots,2N-1$ indexing the measured components of the extended state vector, and $\lfloor\cdot\rfloor$ denoting the floor function. This formulation ensures that both the real and imaginary components are measured with the same error variance, proportional to the energy of the corresponding shell. Notice that the error, chosen to be $5\%$ of the standard deviation of the measured variables, is relatively large when compared to the large-scale variability, where the statistics are approximately Gaussian, and very small at small scales, where energy fluctuations for each shell can easily reach 8-10 times the standard deviation (see Fig. 1(c)).
The diagonal structure of $\mathbf{R}$ reflects the assumption that measurement errors across different shells are independent, implying that the observation process does not introduce cross-scale error correlations. The same structure is used for the perturbation noise required by the EnKF procedure.
@@
[…]
@@
Figure 3: (a) Real parts of the velocity components for shells $n=1$, 66, and $13$ from an experiment where $u_{6}$, $u_{7}$, and $u_{8}$ are measured with observation interval $\Delta t_{\text{obs}}=\tau_{15}=0.002\tau_{0}$. The figure displays the full ensemble (light blue), its mean (solid blue line), and the ground truth (solid black line), over a total experiment duration of $20\tau_{0}$. (b) Probability density function (PDF) of the normalized real-part error, $\frac{\Re(u_{n}-\tilde{u}_{n})^{2}}{\sqrt{\langle\Re(u_{n})^{2}\rangle_{T}}\sqrt{\langle\Re(\tilde{u}_{n})^{2}\rangle_{T}}}$, and (c) PDF of the phase difference, $\theta_{n}-\tilde{\theta}_{n}$, shown for all ensemble members (shaded area) and their mean (blue line) for the selected shells. The PDFs in (b) and (c) are computed by collecting data only after the transient phase—i.e., after saturation is reached (as indicated by the red vertical line in panel (a))—and up to the end of the $20\tau_{0}$ experiment. The light-green shaded zone highlights the time window $12\leq t/\tau_{0}\leq 14$ detailed in Fig. 4(a) and (b).
Figure 4: (a) Ground truth energy spectrum $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$, EnKF estimate $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$, and $\ell^{2}$-error $\mathcal{E}_{n}=\langle|u_{n}-\tilde{u}_{n}|^{2}\rangle_{T,L}$. The grey shaded area highlights the observed shells, while the vertical solid line marks the scale for which the turnover time matches the observation interval, $\tau_{15}=\Delta t_{\textit{obs}}$. (b) Normalized error $\mathcal{E}_{n}/\sqrt{E_{n}\tilde{E}_{n}}$, with the dashed line indicating the scale beyond which statistical consistency ($E_{n}=\tilde{E}_{n}$) is maintained, but the EnKF no longer achieves effective reconstruction. All quantities in panels (a) and (b), including error bars, are averaged over $N_{\text{exp}}=16$ independent experiments, as described in Section IV. (c) Normalized energy evolution for inertial-range scales during the stationary synchronization regime ($12\leq t/\tau_{0}\leq 14$; see light-green shaded region in Fig. 3(a)). Shown are the ground truth values $|u_{n}|^{2}k_{n}^{2/3}$ (left), one randomly selected posterior ensemble member $|\tilde{u}_{n}|^{2}k_{n}^{2/3}$ (middle), and their difference (right). Red vertical lines indicate the measured shells ($n=6$, 77, and 88). (d) Cumulative instantaneous energy flux evolution over the same stationary synchronization regime, shown for the ground truth (left), one randomly selected posterior ensemble member (middle), and their difference (right).
@@ V Ensemble Kalman Filter Results
In our parameter setting, the turnover time of the slowest scale is $\tau_{0}\approx 0.5$, while at the fastest dissipative cutoff scale, it decreases to approximately $\tau_{15}=0.002\tau_{0}=100dt$. Sampling at these respective rates (that is, using $\Delta t_{\textit{obs}}=\tau_{n}$) provides discrete or (quasi-) continuous-time information on inertial range dynamics.
@@
In the first set of experiments, observations are taken from shells $n=6,7,8$, using an observation interval of $\Delta t_{\textit{obs}}=\tau_{15}$ and an ensemble size of $L=1000$ members. Figure 3(a) provides a qualitative overview of the assimilation performance, showing the temporal evolution of the EnKF-based estimates for shells $n=1$, 66, and $13$. The ensemble estimates are continuously updated across all scales, rapidly synchronizing the directly observed $u_{6}$ (middle panel), whose ensemble spread quickly converges within the measurement error bounds. Dynamical correlations also enable full synchronization of distant, unobserved shells such as fast $u_{13}$ and $u_{1}$, albeit with a slight delay. With real-part synchronization already established (Fig. 3(b)), the narrow PDF of $\theta-\tilde{\theta}$ (Fig. 3(c)) completes the reconstruction picture by confirming phase alignment, which in turn implies synchronization of the imaginary components as well.
@@
Figure 6: (a) Ground truth energy $E_{n}=\langle|u_{n}|^{2}\rangle_{T}$ (black dots beneath the colored points) and estimated energy $\tilde{E}_{n}=\langle|\tilde{u}_{n}|^{2}\rangle_{T,L}$ for different assimilation windows $\Delta t_{\textit{obs}}$. The inset shows the corrispondence with scale turnover times $\tau_{n}$. The grey shaded area highlights the measured scales 6, 7, and 8. (b) Synchronization error $\mathcal{E}_{n}$ (Eq.(16)) normalized with energies, for various observation frequencies, with a horizontal dashed line indicating the baseline for statistical independence on non-assimilated scales. (c) Flux-based error $\mathcal{E}^{\Pi}_{n}$ (Eq.(17)) normalized (see Evaluation Metrics for details on normalization). All quantities displayed (including error bars) are averaged over $N_{\text{exp}}=16$ experiments, as described in Evaluation Metrics.
@@
In Fig. 4, we summarize the main quantitative results of this first DA experiment. First, note that the EnKF accurately reproduces the statistical properties of the ground truth, as evidenced in Fig. 4(a) by the perfect overlap between the ground-truth energy spectrum $E_{n}$ and its EnKF-based estimate $\tilde{E}_{n}$.
@@
More interestingly, the instantaneous reconstruction is also reasonably accurate (Fig. 4(a)). This is assessed by comparing the energy spectrum $E_{n}$ with the reconstruction $\ell^{2}$-error $\mathcal{E}_{n}$, as defined in Eq. (16). As shown, the reconstruction is nearly perfect on the observed shells (gray region), and remains highly accurate for all slower shells ($n<6$). For faster scales, good reconstruction is achieved up to $n=16$, where $\mathcal{E}_{n}<E_{n}$, while at higher shell numbers ($n\geq 17$), the reconstruction becomes purely statistical with $\mathcal{E}_{n}\approx 2E_{n}$. The quality of the instantaneous reconstruction is further quantified in Fig. 4(b), where we plot the dimensionless $\ell^{2}$-error $\mathcal{E}_{n}/\sqrt{E_{n}\tilde{E}_{n}}$, normalized by the energy spectra of the ground truth and assimilated fields (Eq. (18)). Complete synchronization with the ground truth, i.e. a practically vanishing normalized error, is achieved up to the fastest observed shell ($n=8$), followed by a partially synchronized range ($9\leq n\leq 16$). Panels (c) and (d) provide a qualitative overview of the time-dependent performance of data assimilation, showing both the normalized energy, $k_{n}^{2/3}|u_{n}|^{2}$, and the energy flux, $\Pi_{n}$ (see Eq. (6)), up to the last inertial shell ($n=13$), over a time window encompassing several energy bursts propagating from large to small scales. The left, central, and right panels, respectively, display the ground-truth quantity, a randomly selected ensemble member, and their pointwise difference. A key observation is that, although the presence of dark halos for $1\leq n\leq 8$ (third column of Panel (c)) confirms the accurate instantaneous tracking of both measured and large-scale dynamics, the most pronounced desynchronization events—evident in the same panel and even more so in the highly intermittent flux shown in Panel (d)—are strongly correlated with intense energy bursts, predicted with high temporal fidelity but limited magnitude accuracy.
@@ V.1 Performance Against Measurements Frequency
The magnitude of the error bars also reflects the sensitivity to the inflation factor (as discussed in Subsection III.1) used for stabilization. For example, the curve corresponding to $\tau_{7}$ in Fig. 6(b) exhibits a pronounced increase in error-bar length, estimated as in Eq. (19), at scales beyond those directly observed, signaling that assimilation was less stable in those cases. This pattern aligns with the theoretical insights of Gottwald and Majda [17], who showed that the stability of the filter critically depends on the observation interval $\Delta t_{\textit{obs}}$. Specifically, when observations are too frequent relative to the relaxation rate toward the attractor, the forecast step is too short for the model dynamics to evolve to overly constrained and potentially unrealistic states. Conversely, when observations are too sparse, the ensemble can spread sufficiently to capture the attractor structure. In other words, while filter divergence is unlikely at either extreme (very small or very large $\Delta t_{\textit{obs}}$) there exists an intermediate regime where the ensemble may be neither sufficiently constrained by observations nor sufficiently dispersed to maintain dynamical consistency, resulting in degraded performance and potential divergence.
The total errors are summarized in Table 1 by summing the errors from the first scale (excluding the forced one) up to the dissipative cutoff at scale 15: $\mathcal{E}=\sum_{n=1}^{15}\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}}$ and $\mathcal{E}^{\Pi}=\sum_{n=1}^{15}\mathcal{E}_{n}^{\Pi}/\sqrt{\tilde{\Delta}_{n}\Delta_{n}}$. A similar procedure is applied to the error bars to obtain the total uncertainties $\Delta\mathcal{E}$ and $\Delta\mathcal{E}^{\Pi}$. Furthermore, the last column of Table 1 reports the number of independent assimilation experiments (out of $N{{}_{\text{exp}}}=16$) that required the correct adjustment of the inflation strength factor $\lambda$, as defined in Eq. (13).
@@
6, 7, 8 $\mathcal{E}\pm\Delta\mathcal{E}$ $\mathcal{E}_{\Pi}\pm\Delta\mathcal{E}_{\Pi}$ $N_{\text{inflation}}$
$\Delta t_{\text{obs}}=\tau_{15}=0.002\tau_{0}$ $4.85\pm 0.26$ $5.11\pm 0.85$ 11
$\Delta t_{\text{obs}}=\tau_{9}=0.02\tau_{0}$ $7.75\pm 0.17$ $8.13\pm 0.98$ 11
$\Delta t_{\text{obs}}=\tau_{8}=0.04\tau_{0}$ $9.10\pm 0.69$ $10.07\pm 3.21$ 66
$\Delta t_{\text{obs}}=\tau_{7}=0.06\tau_{0}$ $14.68\pm 3.90$ $17.76\pm 8.35$ 88
$\Delta t_{\text{obs}}=\tau_{6}=0.1\tau_{0}$ $22.44\pm 1.14$ $27.89\pm 5.20$ 0
$\Delta t_{\text{obs}}=\tau_{4}=0.2\tau_{0}$ $23.73\pm 0.50$ $29.49\pm 4.23$ 0
Table 1: The total error is computed by summing the normalized errors from the first scale $n=1$ (excluding the forced one) up to the dissipative cutoff at scale $n=15$: $\mathcal{E}=\sum_{n=1}^{15}\mathcal{E}_{n}/\sqrt{\tilde{E}_{n}E_{n}}$ and $\mathcal{E}_{\Pi}=\sum_{n=1}^{15}\mathcal{E}_{n}^{\Pi}/\sqrt{\tilde{\Delta}_{n}\Delta_{n}}$. A similar summation is applied to the error bars to obtain $\Delta\mathcal{E}$ and $\Delta\mathcal{E}^{\Pi}$. Details on the number of experiments (out of $N_{\text{exp}}=16$) that required proper tuning of the inflation strength factor are also provided.